solve

xy'+2y=x2lnx


1
Expert's answer
2021-08-09T14:47:21-0400

"xy'+2y=x^2\ln x"

"x^2y'+2xy=x^3\ln x"

"(x^2y)'=x^3\ln x"

"x^2y=\int x^3\ln x dx=\frac{1}{4}\int \ln x dx^4=\frac{x^4}{4}\ln x-\frac{1}{4}\int x^4d\ln x="

"=\frac{x^4}{4}\ln x-\frac{1}{4}\int x^3dx=\frac{x^4}{4}\ln x-\frac{x^4}{16}+C"

"y=\frac{x^2}{16}(4\ln x-1)+Cx^{-2}"